OpenStudy (anonymous):

A high-voltage transmission line with a resistance of 0.315 Ω/km carries a current of 1.05 kA. The line is at a potential of 750 kV at the power station and carries the current to a city located 175 km from the power station. What is the power loss due to resistance in the line?

OpenStudy (anonymous):

I got this to be 6.078×101 MW

OpenStudy (anonymous):

but now it wants to know-->What fraction of the transmitted power does this loss represent?

OpenStudy (anonymous):

Well, we know that \(P=IV\), then, what does this tell us?

OpenStudy (anonymous):

So you have 60780000/750000000?

OpenStudy (anonymous):

Around there, yes.

OpenStudy (anonymous):

it wants it in a simplified fraction,could you help me out on that?

OpenStudy (anonymous):

Well, work it out without solving for variables first. We know that the energy lost is given by \[ P_l=I^2R \]And we know that the total power is: \[ P_T=IR \]Hence: \[ \frac{P_l}{P_T}=\frac{I^2R}{IV} \]Simplify that down and then plug in the numbers.

OpenStudy (anonymous):

Sorry, for the second equation it should be: \[ P_T=IV \]